OpenStudy (anonymous):

find the value of 7^2882 mod 2925 using euler's function i know answer is 49 . but dont know how to figure it out with the euler's function which i need to be able to do

OpenStudy (anonymous):

if i recall correctly \(b^{\phi(m)}\equiv 1(m)\)

OpenStudy (anonymous):

so i guess the first step is to find \(\phi(m)\) which as i recall we do by finding the prime factorization of \(2925\) and then use the fact that \(\phi\) is multiplicative stop me if i am on the wrong track

OpenStudy (anonymous):

yes i think so trying to figure it out going step by step 1 its awkward though !

OpenStudy (anonymous):

\(2925=3^2\times 5^2\times 13\) and so \(\phi(2925)=3\times 2\times 5\times 4\times 12=1440\)

OpenStudy (anonymous):

gotta run, but think that is a good start

OpenStudy (anonymous):

cheers

OpenStudy (anonymous):

Now the 49 is clear right

OpenStudy (anonymous):

2882=2×1440+2

OpenStudy (anonymous):

quick question how did you get the factors eg 3^2 x 5^2 x 13 at the starts ? i know they add up to 2925 but how do you know the factos ?

OpenStudy (anonymous):

i would like to say i factored by hand, but i didn't, i just did this http://www.wolframalpha.com/input/?i=factor+2925 on the other hand you know it is divisible by 25 because it ends in a 25 and you also know it is divisible by 9 because the digits add up to 18, so it is not too hard to factor, since it must contain \(3^2\) and \(5^2\)

OpenStudy (anonymous):

ok cheers

OpenStudy (anonymous):

also is it clear how to find \(\phi(3^2\times 5^2\times 13)\) ?

OpenStudy (anonymous):

and that we know \(7^{\phi(m)}\equiv 1(m)\) making \(7^{2880}\equiv 1(2925)\)

OpenStudy (anonymous):

yesbut what about the 2 ? where does that come into it or how did i get 49